If the dog runs 10m to the north and then turns and runs 20 m to the west and traveled for 10 seconds, what is its speed in going from its starting position to its final position. Explain you answer
1
Expert's answer
2019-08-20T10:24:59-0400

According Pythagorean theorem:


"s=\sqrt{10^2+20^2}=10\sqrt{5}m"

Its speed in going from its starting position to its final position:


"v=\frac{s}{t}=\frac{10\sqrt{5}}{10}=\sqrt{5}\frac{m}{s}=2.2\frac{m}{s}"